OpenStudy (anonymous):

what is 7 square root of 20 simplified to?

OpenStudy (anonymous):

\[7\sqrt{20}\] Remember how you can factor under the radical sign... so... \[7\sqrt{20} = 7\sqrt{4\times5} = 7\sqrt{2\times2\times5}\] You can then take out the square root of 4, pulling out the twos to get... \[7\sqrt{20} = 7\times2\sqrt{5}\] Which brings you to: \[14\sqrt{5}\]